# More on soundness in the enriched context

### v2

@@ 1. Introduction
Here we deal with the more general notion of locally $\Phi$-presentable ${\mathcal{V}}$-category, which specialize to that of Kelly when $\Phi$ is the class of finite weighted limits. This was first considered in the ordinary setting in [1] and [16], was partially extended to enriched categories in [29], and then again in a formal 2-categorical framework [18]. With our approach we do not need to assume the base of enrichment to be locally presentable (as in [25]), but merely complete and cocomplete (as in [24, 29]); this allows us to capture some topological examples such as the cartesian closed category $\mathbf{CGTop}$ compactly generated topological spaces.
@@
Above, a ${\mathcal{V}}$-category is called $\Phi$-accessible if it is the free cocompletion of a small ${\mathcal{V}}$-category under $\Phi$-flat colimits [1, 32]. With the same hypotheses on ${\mathcal{V}}$ we can also prove a generalization of the Gabriel–Ulmer duality [21, 25]:
@@
[…]
@@
Ordinarily, and for $\Phi$ and $\Psi$ consisting respectively on the classes of $\alpha$-small and $\beta$-small diagrams (for regular cardinals $\alpha<\beta$), we know that the above condition holds if and only if $\alpha$ is sharply less than $\beta$ in the sense of [36, Definition 2.3.1]. In our more general framework we introduce a notion of sharply less than relationship for sound classes of weights (which restricts to the original one when considering classes induced by regular cardinals) defined by the equivalent conditions below. This generalizes to the enriched settings parts of [33, Theorem 2.3].
@@
    For any small ${\mathcal{V}}$-category ${\mathcal{C}}$ we have an equivalence
    $\Psi\negthinspace^{{}^{+}}\negthinspace(\Phi^{{}^{+}}_{\Psi}{\mathcal{C}})\simeq\Phi\negthinspace^{{}^{+}}\negthinspace{\mathcal{C}}.$
@@
    Every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\Psi$-accessible and there is an equivalence
    $\Phi^{{}^{+}}_{\Psi}{\mathcal{A}}_{\Phi}\simeq{\mathcal{A}}_{\Psi}.$
@@
Part 3. The third and final part of the paper (Section 5) is devoted to the study of those enriched equational theories from [40], which are presented by $\Phi$-ary equations over a $\Phi$-ary language, for a given weakly sound class $\Phi$.
A language in the sense of [40] is a set of function symbols with specified domain and codomain arities $(X,Y)$ varying among the objects of ${\mathcal{V}}$; for each such ${\mathbb{L}}$ there is a corresponding notion of ${\mathbb{L}}$-structure and of ${\mathcal{V}}$-category of ${\mathbb{L}}$-structures. The class of ${\mathbb{L}}$-terms is built recursively starting from the function symbols in ${\mathbb{L}}$ and applying the rules of Definition 5.3. Then, a theory ${\mathbb{E}}$ over ${\mathbb{L}}$ is a set of equations $(s=t)$ involving terms of the same arity; a model of such a theory is an ${\mathbb{L}}$-structure which satisfies these equations, once interpreted. All this background material will be covered in Section 5.1.
Starting from this, we define a ${\mathbb{L}}$ and and ${\mathbb{L}}$-theory ${\mathbb{E}}$ to be $\Phi$-ary if they involve function symbols and terms whose arities are (certain) $\Phi$-presentable objects (see Definition 5.10 for the details). Among others, these $\Phi$-theories are important to characterize ${\mathcal{V}}$-categories of algebras of those enriched monads on ${\mathcal{V}}$ which preserve $\Phi$-flat colimits:
Theorem 5.13.
@@
    ${\mathcal{K}}\simeq\operatorname{\textnormal{M}od}(\mathbb{E})$ for a $\Phi$-ary equational theory $\mathbb{E}$;
@@
In (4) above, $\Phi{\mathcal{I}}$ is the closure of the unit $I$ in ${\mathcal{V}}$ under $\Phi$-colimits. When ${\mathcal{V}}$ is locally $\lambda$-presentable as a closed category and $\Phi$ is the class of the $\lambda$-small weights, this specializes to [40, Theorem 5.14]. In Section 5.3.1 we take ${\mathcal{V}}$ to be cartesian closed and $\Phi$ to consist of the class of finite products; in this case $\Phi$-flat colimits are called ${\mathcal{V}}$-sifted. The theorem above then provides a characterization of the ${\mathcal{V}}$-categories of algebras of strongly finitary (enriched) monads on ${\mathcal{V}}$; these have been studied for instance in [2, 23, 3, 39]. In Appendix A, following [14], we characterize 2-dimensional sifted colimits.
@@ 2. Background
here ${\mathcal{V}}\textnormal{-}\mathbf{CAT}({\mathcal{C}},{\mathcal{K}})$ is the category of all ${\mathcal{V}}$-functors from ${\mathcal{C}}\to{\mathcal{K}}$ and ${\mathcal{V}}$-natural transformations between them, while we denote by $\Phi\textnormal{-Coct}(\Phi{\mathcal{C}},{\mathcal{K}})$ the full subcategory of ${\mathcal{V}}\textnormal{-}\mathbf{CAT}(\Phi{\mathcal{C}},{\mathcal{K}})$ spanned by the $\Phi$-cocontinuous ${\mathcal{V}}$-functors. The inverse of the equivalence above is given by left Kan extending along $Z$ (see [27, Proposition 3.6]).
@@
If $\Phi$ is a small set of weights, then $\Phi$ is locally small. The class ${\mathcal{Q}}$ of the Cauchy weights is locally small whenever ${\mathcal{V}}$ is locally presentable (Corollary 3.33 or [22]). The class ${\mathcal{P}}$ of all small weights is not locally small in general.
@@
Every saturated class is pre-saturated by [8]. Pre-saturated classes have been studied for instance in [42, Appendix A.1]. Note that most commonly used classes of weights are pre-saturated, but in general not saturated. Indeed, the class of finite limits is pre-saturated but not saturated; its saturation is the class of L-finite limits (see [38, Section 3]).
@@
A locally small class $\Phi$ is instead called sound [32, Definition 3.2] if for any weight $M$, whenever $M*-$ preserves $\Phi$-limits of representables, then $M$ is $\Phi$-flat. This corresponds to the notion of soundness originally introduced in [1], but is somewhat harder to deal with.
@@
See [33, Example 4.8] for details. In cases where ${\mathcal{V}}=\mathbf{Set}$ and we speak of a class of categories rather than a class of weights, we are referring to the corresponding conical limits.
@@
    ${\mathcal{V}}$ cartesian closed, $\Phi$ is the class for finite products. Examples of such ${\mathcal{V}}$ are the categories $\mathbf{Pos}$ of partially ordered sets, $\mathbf{Cat}$ of small categories, and $\mathbf{SSet}$ of simplicial sets; but also ${\mathcal{V}}=\mathbf{CGTop},\mathbf{HCGTop}$, and $\mathbf{QTop}$ of compactly generated topological spaces, Hausdorff such, and quasi-topological spaces, which are (locally bounded but) not locally presentable. ([23, Lemma 2.3]).
@@
We can characterize soundness as follows:
@@
with $W\colon\Phi{\mathcal{C}}\hookrightarrow[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]$. Since $\Phi{\mathcal{C}}\subseteq[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]_{\Phi}$ (essentially by definition) and $\textnormal{Ran}_{J^{\textnormal{op}}}M$ is $\Phi$-flat, one concludes thanks to [33, 4.9] which says that $\Phi$ is weakly sound if and only if every object of $[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]$ is a $\Phi$-flat colimit of objects from $\Phi{\mathcal{C}}$. ∎
@@ 3.1. Definitions and basic results
The following was first considered in the enriched context in [29, Section 6.3] with an equivalent definition. They say that a ${\mathcal{V}}$-category is locally $\Phi$-presentable if it is $\Phi$-accessible (Definition 3.5) and cocomplete, while we follow the approach that involves a strong generator. That these provide equivalent notions will follow from our Theorem 3.9.
@@
The notion of $\Phi$-accessible ${\mathcal{V}}$-category has been already introduced in [32, Section 3] for a locally presentable base of enrichment. They can be defined exactly in the same way also in this context (see below), and most of the results of [32] extend with the only change that Cauchy completions may not be small.
@@ 3.2. Characterization theorems
    if ${\mathcal{C}}$ is the closure of ${\mathcal{G}}$ under $\Phi$-colimits, then ${\mathcal{C}}$ is small, dense, its Cauchy completion is ${\mathcal{K}}_{\Phi}$, and every $X\in{\mathcal{K}}$ can be expressed as the $\Phi$-flat colimit:
    $X\cong{\mathcal{K}}(H_{\mathcal{C}}-,X)*H_{\mathcal{C}},$
    where $H_{\mathcal{C}}\colon{\mathcal{C}}\to{\mathcal{K}}$ is the inclusion.
@@
    every $X\in{\mathcal{K}}$ can be expressed as the (possibly large) $\Phi$-flat colimit:
    $X\cong{\mathcal{K}}(H-,X)*H.$
@@
The points (1, 2, 4) are proved as in [32, Proposition 3.9] with the only difference that one should keep in mind that the Cauchy completion ${\mathcal{Q}}({\mathcal{C}})$ of a small ${\mathcal{V}}$-category ${\mathcal{C}}$ may not be small, and that for any $X\in{\mathcal{V}}$ the restriction ${\mathcal{K}}(J-,X)\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ is $\Phi$-flat (the inclusion $J\colon{\mathcal{C}}\to{\mathcal{K}}$ is $\Phi$-cocontinuous).
@@
Density then implies that $X\cong{\mathcal{K}}(H_{\mathcal{C}}-,X)*H_{\mathcal{C}}$ for any $X\in{\mathcal{K}}$; this colimit is $\Phi$-flat since ${\mathcal{K}}(H_{\mathcal{C}}-,X)$ is $\Phi$-continuous. The fact that ${\mathcal{K}}_{\Phi}$ is the Cauchy completion of ${\mathcal{C}}$ follows from the same arguments as [32, Proposition 3.9]. ∎
@@
$(1)\Rightarrow(2)$. Let $F\colon{\mathcal{K}}\to{\mathcal{A}}$ be continuous and $\Phi$-flat colimit preserving. We need to prove that for each $A\in{\mathcal{A}}$ the ${\mathcal{V}}$-functor ${\mathcal{A}}(A,F-)$ is representable. Since ${\mathcal{A}}$ is $\Phi$-accessible it is enough to consider only the $\Phi$-presentable $A\in{\mathcal{A}}$ (then the left adjoint extends under $\Phi$-flat colimits to all of ${\mathcal{A}}$). Consider then $A\in{\mathcal{A}}_{\Phi}$; by [24, Theorem 4.80] to prove that ${\mathcal{A}}(A,F-)$ is representable is the same as proving that $\{{\mathcal{A}}(A,F-),\textnormal{id}_{{\mathcal{K}}}\}$ exists and is preserved by ${\mathcal{A}}(A,F-)$. By hypothesis ${\mathcal{A}}(A,F-)$ preserves $\Phi$-flat colimits and ${\mathcal{K}}$ is $\Phi$-accessible; thus ${\mathcal{A}}(A,F-)$ is the left Kan extension of its restriction to the $\Phi$-presentable objects. Let $H\colon{\mathcal{K}}_{\Phi}\to{\mathcal{K}}$ be the inclusion; then this means that ${\mathcal{A}}(A,F-)\cong{\mathcal{A}}(A,FH-)*YH^{\textnormal{op}}$, where $Y\colon{\mathcal{K}}^{\textnormal{op}}\to[{\mathcal{K}},{\mathcal{V}}]$ is Yoneda. We can consider then the following chain of isomorphisms, each side existing if the other does:
$\begin{split}\{{\mathcal{A}}(A,F-),\textnormal{id}_{{\mathcal{K}}}\}&\cong\{{\mathcal{A}}(A,FH-)*YH^{\textnormal{op}},\textnormal{id}_{{\mathcal{K}}}\}\\ &\cong\{{\mathcal{A}}(A,FH-),\{YH^{\textnormal{op}},\textnormal{id}_{{\mathcal{K}}}\}\}\\ &\cong\{{\mathcal{A}}(A,FH-),H^{\textnormal{op}}\}\\ \end{split}$
@@
[…]
@@
The equivalence $(2)\Leftrightarrow(4)$ below is given by [29, 6.4]. While, when ${\mathcal{V}}_{0}$ is locally presentable, the first three equivalences and the fifth follow from [33, 4.22].
@@
Let $\mathbf{L}\Phi\mathbf{P}$ be the 2-category of locally $\Phi$-presentable ${\mathcal{V}}$-categories, continuous and $\Phi$-flat colimit preserving ${\mathcal{V}}$-functors, and ${\mathcal{V}}$-natural transformations. On the other side let $\Phi\textnormal{-}\mathbf{Cont}$ be the 2-category of $\Phi$-complete ${\mathcal{V}}$-categories which are the Cauchy completion of a small ${\mathcal{V}}$-category, $\Phi$-continuous ${\mathcal{V}}$-functors, and ${\mathcal{V}}$-natural transformation.
@@
[…]
@@
Let ${\mathcal{K}}$ be locally $\Phi$-presentable; then we first note that $\mathbf{L}\Phi\mathbf{P}({\mathcal{K}},{\mathcal{V}})\simeq{\mathcal{K}}_{\Phi}^{\textnormal{op}}$, induced by the restriction of the Yoneda embedding $Y\colon{\mathcal{K}}^{\textnormal{op}}\to[{\mathcal{K}},{\mathcal{V}}]$. It is clear that for every $X\in{\mathcal{K}}_{\Phi}$ we have ${\mathcal{K}}(X,-)\in\mathbf{L}\Phi\mathbf{P}({\mathcal{K}},{\mathcal{V}})$. Conversely let $F\colon{\mathcal{K}}\to{\mathcal{V}}$ be continuous and $\Phi$-flat colimits preserving; then by the previous proposition it has a left adjoint $L$. As a consequence ${\mathcal{K}}(LI,-)\cong[I,F-]\cong F$ and $LI\in{\mathcal{K}}_{\Phi}$, proving the equality. Thus the 2-functor is well defined since ${\mathcal{K}}_{\Phi}^{\textnormal{op}}$ is $\Phi$-complete and the Cauchy completion of a small category. Moreover, given $F\colon{\mathcal{K}}\to{\mathcal{A}}$ in $\mathbf{L}\Phi\mathbf{P}$ it follows that $\mathbf{L}\Phi\mathbf{P}(F-,{\mathcal{V}})$ corresponds, up to the equivalence above, to the restriction of the left adjoint of $F$ to the $\Phi$-presentable objects. Now, the fact that $\Phi\textnormal{-}\mathbf{Cont}(\mathbf{L}\Phi\mathbf{P}({\mathcal{K}},{\mathcal{V}}),{\mathcal{V}})\cong{\mathcal{K}}$ follows directly from $(1)\Rightarrow(2)$ of Theorem 3.9.

### v3

@@ 1. Introduction
Here we deal with the more general notion of locally $\Phi$-presentable ${\mathcal{V}}$-category, which specializes to that of Kelly when $\Phi$ is the class of finite weighted limits. This was first considered in the ordinary setting in [1] and [16], was partially extended to enriched categories in [30], and then again in a formal 2-categorical framework [18]. With our approach we do not need to assume the base of enrichment to be locally presentable (as in [26]), but merely complete and cocomplete (as in [25, 30]); this allows us to capture some topological examples such as the cartesian closed category $\mathbf{CGTop}$ compactly generated topological spaces.
@@
Above, a ${\mathcal{V}}$-category is called $\Phi$-accessible if it is the free cocompletion of a small ${\mathcal{V}}$-category under $\Phi$-flat colimits [1, 33]. With the same hypotheses on ${\mathcal{V}}$ we can also prove a generalization of the Gabriel–Ulmer duality [22, 26, 16]:
@@
$\mathbf{L}\Phi\mathbf{P}(-,{\mathcal{V}})\colon\mathbf{L}\Phi\mathbf{P}$$\Phi\textnormal{-}\mathbf{Cont}^{\textnormal{op}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0mu\Phi\textnormal{-}\mathbf{Cont}(-,{\mathcal{V}})$
@@
Ordinarily, and for $\Phi$ and $\Psi$ consisting respectively on the classes of $\alpha$-small and $\beta$-small diagrams (for regular cardinals $\alpha<\beta$), we know that the above condition holds if and only if $\alpha$ is sharply less than $\beta$ in the sense of [38, Definition 2.3.1]. In our more general framework we introduce a notion of sharply less than relationship for sound classes of weights (which restricts to the original one when considering classes induced by regular cardinals) defined by the equivalent conditions below. This generalizes to the enriched settings parts of [34, Theorem 2.3].
@@
    For any small ${\mathcal{V}}$-category ${\mathcal{C}}$ we have an equivalence Ψ^^+(Φ^^+_ΨC) Φ^^+C.
@@
    Every $\Phi$-accessible ${\mathcal{V}}$-category ${\mathcal{A}}$ is $\Psi$-accessible and there is an equivalence Φ^^+_ΨAA_Ψ.
@@
Part 3. The third and final part of the paper (Section 5) is devoted to the study of those enriched equational theories from [42], which are presented by $\Phi$-ary equations over a $\Phi$-ary language, for a given weakly sound class $\Phi$. As for [42], the point of this Section is to show that equations can be defined over a class of recursively generated terms, thus moving in a different direction with respect to earlier works [36, 21].
A language in the sense of [36, 42] is a set of function symbols with specified domain and codomain arities $(X,Y)$ varying among the objects of ${\mathcal{V}}$; for each such ${\mathbb{L}}$ there is a corresponding notion of ${\mathbb{L}}$-structure and of ${\mathcal{V}}$-category of ${\mathbb{L}}$-structures. The class of ${\mathbb{L}}$-terms is built recursively starting from the function symbols in ${\mathbb{L}}$ and applying the rules of Definition 5.3 (and is a subclass of the terms considered in [36]). Then, a theory ${\mathbb{E}}$ over ${\mathbb{L}}$ is a set of equations $(s=t)$ involving terms of the same arity; a model of such a theory is an ${\mathbb{L}}$-structure which satisfies these equations, once interpreted. All this background material will be covered in Section 5.1.
Starting from this, we define a language ${\mathbb{L}}$ and an ${\mathbb{L}}$-theory ${\mathbb{E}}$ to be $\Phi$-ary if they involve function symbols and terms whose arities are (certain) $\Phi$-presentable objects (see Definition 5.10 for the details). Among others, $\Phi$-theories characterize ${\mathcal{V}}$-categories of algebras of those enriched monads on ${\mathcal{V}}$ which preserve $\Phi$-flat colimits:
Theorem (part of 5.13).
@@
    ${\mathcal{K}}\simeq\operatorname{\textnormal{M}od}(\mathbb{E})$ for a $\Phi$-ary equational theory $\mathbb{E}$ with recursively generated terms;
@@
In (4) above, $\Phi{\mathcal{I}}$ is the closure of the unit $I$ in ${\mathcal{V}}$ under $\Phi$-colimits. When ${\mathcal{V}}$ is locally $\lambda$-presentable as a closed category and $\Phi$ is the class of the $\lambda$-small weights, this specializes to [42, Theorem 5.14]. In Section 5.3.1 we take ${\mathcal{V}}$ to be cartesian closed and $\Phi$ to consist of the class of finite products; in this case $\Phi$-flat colimits are called ${\mathcal{V}}$-sifted. The theorem above then provides a characterization of the ${\mathcal{V}}$-categories of algebras of strongly finitary (enriched) monads on ${\mathcal{V}}$; these have been studied for instance in [2, 24, 3, 41]. In Appendix A, following [14], we characterize 2-dimensional sifted colimits.
@@ 2. Background
here ${\mathcal{V}}\textnormal{-}\mathbf{CAT}({\mathcal{C}},{\mathcal{K}})$ is the category of all ${\mathcal{V}}$-functors from ${\mathcal{C}}\to{\mathcal{K}}$ and ${\mathcal{V}}$-natural transformations between them, while we denote by $\Phi\textnormal{-Coct}(\Phi{\mathcal{C}},{\mathcal{K}})$ the full subcategory of ${\mathcal{V}}\textnormal{-}\mathbf{CAT}(\Phi{\mathcal{C}},{\mathcal{K}})$ spanned by the $\Phi$-cocontinuous ${\mathcal{V}}$-functors. The inverse of the equivalence above is given by left Kan extending along $Z$ (see [28, Proposition 3.6]).
@@
If $\Phi$ is a small set of weights, then $\Phi$ is locally small. The class ${\mathcal{Q}}$ of the Cauchy weights is locally small whenever ${\mathcal{V}}$ is locally presentable ([23] or Corollary 3.35). The class ${\mathcal{P}}$ of all small weights is not locally small in general.
@@
Every saturated class is pre-saturated by [8]. Pre-saturated classes have been studied for instance in [44, Appendix A.1]. Note that most commonly used classes of weights are pre-saturated, but in general not saturated. Indeed, the class of finite limits is pre-saturated but not saturated; its saturation is the class of L-finite limits (see [40, Section 3]).
@@
A locally small class $\Phi$ is instead called sound [33, Definition 3.2] if for any weight $M$, whenever $M*$- preserves $\Phi$-limits of representables, then $M$ is $\Phi$-flat. This corresponds to the notion of soundness originally introduced in [1], but is somewhat harder to deal with.
@@
See [34, Example 4.8] for details. In cases where ${\mathcal{V}}=\mathbf{Set}$ and we speak of a class of categories rather than a class of weights, we are referring to the corresponding conical limits.
@@
    ${\mathcal{V}}$ cartesian closed, $\Phi$ is the class for finite products. Examples of such ${\mathcal{V}}$ are the categories $\mathbf{Pos}$ of partially ordered sets, $\mathbf{Cat}$ of small categories, and $\mathbf{SSet}$ of simplicial sets; but also ${\mathcal{V}}=\mathbf{CGTop},\mathbf{HCGTop}$, and $\mathbf{QTop}$ of compactly generated topological spaces, Hausdorff such, and quasi-topological spaces, which are (locally bounded but) not locally presentable. ([24, Lemma 2.3]).
@@
We can characterize soundness as follows, somewhat expanding the equivalent conditions given by [34, Propositions 4.5 & 4.9].
@@
with $W\colon\Phi{\mathcal{C}}\hookrightarrow[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]$. Since $\Phi{\mathcal{C}}\subseteq[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]_{\Phi}$ (essentially by definition) and $\textnormal{Ran}_{J^{\textnormal{op}}}M$ is $\Phi$-flat, one concludes thanks to [34, 4.9] which says that $\Phi$ is weakly sound if and only if every object of $[{\mathcal{C}}^{\textnormal{op}},{\mathcal{V}}]$ is a $\Phi$-flat colimit of objects from $\Phi{\mathcal{C}}$. ∎
@@ 3.1. Definitions and basic results
The following was first considered in the enriched context in [30, Section 6.3] with an equivalent definition. They say that a ${\mathcal{V}}$-category is locally $\Phi$-presentable if it is $\Phi$-accessible (Definition 3.5) and cocomplete, while we follow the approach that involves a strong generator. That these provide equivalent notions will follow from our Theorem 3.9.
@@
The notion of $\Phi$-accessible ${\mathcal{V}}$-category has been already introduced in [33, Section 3] for a locally presentable base of enrichment. They can be defined exactly in the same way also in this context (see below), and most of the results of [33] extend with the only change that Cauchy completions may not be small.
@@ 3.2. Characterization theorems
    if ${\mathcal{C}}$ is the closure of ${\mathcal{G}}$ under $\Phi$-colimits, then ${\mathcal{C}}$ is small, dense, its Cauchy completion is ${\mathcal{K}}_{\Phi}$, and every $X\in{\mathcal{K}}$ can be expressed as the $\Phi$-flat colimit: XK(H_C-,X)*H_C, where $H_{\mathcal{C}}\colon{\mathcal{C}}\to{\mathcal{K}}$ is the inclusion.
@@
    every $X\in{\mathcal{K}}$ can be expressed as the (possibly large) $\Phi$-flat colimit: XK(H-,X)*H.
@@
The points (1, 2, 4) are proved as in [33, Proposition 3.9] with the only difference that one should keep in mind that the Cauchy completion ${\mathcal{Q}}({\mathcal{C}})$ of a small ${\mathcal{V}}$-category ${\mathcal{C}}$ may not be small, and that for any $X\in{\mathcal{V}}$ the restriction ${\mathcal{K}}(J-,X)\colon{\mathcal{C}}^{\textnormal{op}}\to{\mathcal{V}}$ is $\Phi$-flat (the inclusion $J\colon{\mathcal{C}}\to{\mathcal{K}}$ is $\Phi$-cocontinuous).
@@
Density then implies that $X\cong{\mathcal{K}}(H_{\mathcal{C}}-,X)*H_{\mathcal{C}}$ for any $X\in{\mathcal{K}}$; this colimit is $\Phi$-flat since ${\mathcal{K}}(H_{\mathcal{C}}-,X)$ is $\Phi$-continuous. The fact that ${\mathcal{K}}_{\Phi}$ is the Cauchy completion of ${\mathcal{C}}$ follows from the same arguments as [33, Proposition 3.9]. ∎
@@
$(1)\Rightarrow(2)$. Let $F\colon{\mathcal{K}}\to{\mathcal{A}}$ be continuous and $\Phi$-flat colimit preserving. We need to prove that for each $A\in{\mathcal{A}}$ the ${\mathcal{V}}$-functor ${\mathcal{A}}(A,F-)$ is representable. Since ${\mathcal{A}}$ is $\Phi$-accessible it is enough to consider only the $\Phi$-presentable $A\in{\mathcal{A}}$ (then the left adjoint extends under $\Phi$-flat colimits to all of ${\mathcal{A}}$). Consider then $A\in{\mathcal{A}}_{\Phi}$; by [25, Theorem 4.80] to prove that ${\mathcal{A}}(A,F-)$ is representable is the same as proving that $\{{\mathcal{A}}(A,F-),\textnormal{id}_{{\mathcal{K}}}\}$ exists and is preserved by ${\mathcal{A}}(A,F-)$. By hypothesis ${\mathcal{A}}(A,F-)$ preserves $\Phi$-flat colimits and ${\mathcal{K}}$ is $\Phi$-accessible; thus ${\mathcal{A}}(A,F-)$ is the left Kan extension of its restriction to the $\Phi$-presentable objects. Let $H\colon{\mathcal{K}}_{\Phi}\to{\mathcal{K}}$ be the inclusion; then this means that ${\mathcal{A}}(A,F-)\cong{\mathcal{A}}(A,FH-)*YH^{\textnormal{op}}$, where $Y\colon{\mathcal{K}}^{\textnormal{op}}\to[{\mathcal{K}},{\mathcal{V}}]$ is Yoneda. We can consider then the following chain of isomorphisms, each side existing if the other does:
[…]
@@
$(2)\Rightarrow(3)$. Let $L$ be the left adjoint to $F$ and $A\in{\mathcal{A}}_{\Phi}$; then ${\mathcal{K}}(LA,-)\cong{\mathcal{A}}(A,F-)$ preserves $\Phi$-flat colimits since $F$ and ${\mathcal{A}}(A,-)$ do.
@@
The equivalence $(2)\Leftrightarrow(4)$ below is given by [30, 6.4]. While, when ${\mathcal{V}}_{0}$ is locally presentable, the first three equivalences and the fifth follow from [34, 4.22].
@@
Example 3.11.
Consider a cartesian closed base of enrichment ${\mathcal{V}}$ together with the sound class $\mathbf{Fp}$ for finite products (Example 2.6). Given an ordinary category ${\mathcal{C}}$ with finite products, it is easy to see that the free ${\mathcal{V}}$-category ${\mathcal{C}}_{\mathcal{V}}$ over ${\mathcal{C}}$ still has finite products and that these are computed as in ${\mathcal{C}}$. It follows that the ${\mathcal{V}}$-category $\textnormal{Fp}({\mathcal{C}}_{\mathcal{V}},{\mathcal{V}})$ is locally Fp-presentable (by Theorem 3.9), so that all the results of this section apply. But if we take its underlying category, by construction we have that
$\textnormal{Fp}({\mathcal{C}}_{\mathcal{V}},{\mathcal{V}})_{0}\cong\textnormal{Fp}({\mathcal{C}},{\mathcal{V}}_{0})$
which is in general not locally $\mathbf{Fp}$-presentable (that is, not a finitary variety), nor locally presentable unless ${\mathcal{V}}_{0}$ is.
For instance, we can consider ${\mathcal{V}}=\mathbf{CGTop}$, the cartesian closed category of compactly generated topological spaces (which is not a variety and not locally presentable), and the ordinary Lawvere theory for groups. It follows from the arguments above, that the enriched category of groups internal to $\mathbf{CGTop}$ is locally $\mathbf{Fp}$-presentable, but its underlying ordinary category is not.
We now prove an enriched analogue of the Gabriel–Ulmer duality for sound classes that will generalize the ordinary version of [16] (see also [45, Remark 3.8]). Denote by $\mathbf{L}\Phi\mathbf{P}$ the 2-category of locally $\Phi$-presentable ${\mathcal{V}}$-categories, continuous and $\Phi$-flat colimit preserving ${\mathcal{V}}$-functors, and ${\mathcal{V}}$-natural transformations. On the other hand, let $\Phi\textnormal{-}\mathbf{Cont}$ be the 2-category of $\Phi$-complete ${\mathcal{V}}$-categories which are the Cauchy completion of a small ${\mathcal{V}}$-category, $\Phi$-continuous ${\mathcal{V}}$-functors, and ${\mathcal{V}}$-natural transformation.
Notation 3.12.
Given ${\mathcal{K}}\in\mathbf{L}\Phi\mathbf{P}$, we denote with $\mathbf{L}\Phi\mathbf{P}({\mathcal{K}},{\mathcal{V}})$ — bolded — the hom-category in $\mathbf{L}\Phi\mathbf{P}$. This is the (ordinary) category of continuous and $\Phi$-flat colimit preserving ${\mathcal{V}}$-functors ${\mathcal{K}}\to{\mathcal{V}}$, and ${\mathcal{V}}$-natural transformations between them. Instead, we denote by $\textnormal{L}\Phi\textnormal{P}({\mathcal{K}},{\mathcal{V}})$ — unbolded — the ${\mathcal{V}}$-category obtained by taking the full subcategory of $[{\mathcal{K}},{\mathcal{V}}]$ spanned by the continuous and $\Phi$-flat colimit preserving ${\mathcal{V}}$-functors (the fact that this is actually a ${\mathcal{V}}$-category, and not just a ${\mathcal{V}}^{\prime}$-category for some universe enlargement ${\mathcal{V}}^{\prime}\supseteq{\mathcal{V}}$, follows from the theorem below). Therefore we obtain an equality
$\textnormal{L}\Phi\textnormal{P}({\mathcal{K}},{\mathcal{V}})_{0}=\mathbf{L}\Phi\mathbf{P}({\mathcal{K}},{\mathcal{V}}).$
We use the same notational convention for ${\mathcal{C}}\in\Phi\textnormal{-}\mathbf{Cont}$, where the ${\mathcal{V}}$-category $\Phi\textnormal{-}\textnormal{Cont}({\mathcal{C}},{\mathcal{V}})$ has already been defined. Therefore we obtain an equality
$\Phi\textnormal{-}\textnormal{Cont}({\mathcal{C}},{\mathcal{V}})_{0}=\Phi\textnormal{-}\mathbf{Cont}({\mathcal{C}},{\mathcal{V}})$
of ordinary categories.
@@
$\textnormal{L}\Phi\textnormal{P}(-,{\mathcal{V}})\colon\mathbf{L}\Phi\mathbf{P}$$\Phi\textnormal{-}\mathbf{Cont}^{\textnormal{op}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0mu\Phi\textnormal{-}\textnormal{Cont}(-,{\mathcal{V}})$
@@
Let ${\mathcal{K}}$ be locally $\Phi$-presentable; then we first note that $\textnormal{L}\Phi\textnormal{P}({\mathcal{K}},{\mathcal{V}})\simeq{\mathcal{K}}_{\Phi}^{\textnormal{op}}$, induced by the restriction of the Yoneda embedding $Y\colon{\mathcal{K}}^{\textnormal{op}}\to[{\mathcal{K}},{\mathcal{V}}]$. It is clear that for every $X\in{\mathcal{K}}_{\Phi}$ we have ${\mathcal{K}}(X,-)\in\textnormal{L}\Phi\textnormal{P}({\mathcal{K}},{\mathcal{V}})$. Conversely let $F\colon{\mathcal{K}}\to{\mathcal{V}}$ be continuous and $\Phi$-flat colimits preserving; then by the previous proposition it has a left adjoint $L$. As a consequence ${\mathcal{K}}(LI,-)\cong[I,F-]\cong F$ and $LI\in{\mathcal{K}}_{\Phi}$, proving the equivalence. Thus the 2-functor is well defined since ${\mathcal{K}}_{\Phi}^{\textnormal{op}}$ is $\Phi$-complete and the Cauchy completion of a small category. Moreover, given $F\colon{\mathcal{K}}\to{\mathcal{A}}$ in $\mathbf{L}\Phi\mathbf{P}$ it follows that $\textnormal{L}\Phi\textnormal{P}(F-,{\mathcal{V}})$ corresponds, up to the equivalence above, to the restriction of the left adjoint of $F$ to the $\Phi$-presentable objects. Now, the fact that $\Phi\textnormal{-}\textnormal{Cont}(\textnormal{L}\Phi\textnormal{P}({\mathcal{K}},{\mathcal{V}}),{\mathcal{V}})\cong{\mathcal{K}}$ follows directly from $(1)\Rightarrow(2)$ of Theorem 3.9.